OpenStudy (anonymous):

\sqrt{a}\ • \sqrt[3]{a^2}\

OpenStudy (jdoe0001):

\(\bf \Large{\sqrt{a} \cdot \sqrt[3]{a^2}\qquad \textit{recall that}\quad \sqrt[m]{a^n}=a^{\frac{n}{m}}\qquad thus\\ \quad \\ \sqrt{a} \cdot \sqrt[3]{a^2}\implies a^{\frac{1}{2}}\cdot a^{\frac{2}{3}}}\)